OpenStudy (anonymous):

here are some equations sets using the digits: 1 to 4 Each digit must be once. can you find some more?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

this is my examples :3-1=4/2

OpenStudy (anonymous):

2-1=4-3 2+3=4+1

OpenStudy (anonymous):

thats already been an examples

OpenStudy (anonymous):

can you give other

OpenStudy (anonymous):

no, not without another digit

OpenStudy (anonymous):

\[1^{2} = 4-3\]

OpenStudy (anonymous):

\[\sqrt[2]{4} = 3-1\]

OpenStudy (anonymous):

i can but it only few,i need another... my examples is: 3-1=4/2 :4-2=3-1

OpenStudy (anonymous):

\[2^{(3-1)} = 4\]

OpenStudy (anonymous):

4 x 3 = 12? lol

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

any?

OpenStudy (anonymous):

i just found em all

OpenStudy (anonymous):

Nah, there's more. 3-4+2=1 for example. Just gotta get creative now.